OpenStudy (anonymous):

Verify each identity: cos^2x-sin^2x=2cos^2x-1

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta)}}\\ \quad \\ \quad \\ cos^2(x)-sin^2(x)\implies cos^2(x)-[{\color{blue}{ 1-cos^2(\theta)}}]\) expand that, see what you get

OpenStudy (anonymous):

Thank you!! :)